### v2

# A Particle Algorithm for Mean-Field Variational Inference

Abstract
Variational inference is a fast and scalable alternative to Markov chain Monte Carlo and has been widely applied to posterior inference tasks in statistics and machine learning. A traditional approach for implementing mean-field variational inference (MFVI) is coordinate ascent variational inference (CAVI), which relies crucially on parametric assumptions on complete conditionals. In this paper, we introduce a novel particle-based algorithm for mean-field variational inference, which we term PArticle VI (PAVI). Notably, our algorithm does not rely on parametric assumptions on complete conditionals, and it applies to the nonparametric setting. We provide non-asymptotic finite-particle convergence guarantee for our algorithm. To our knowledge, this is the first end-to-end guarantee for particle-based MFVI.
Keywords: Variational inference, Wasserstein gradient flow, particle algorithm, mean-field approximation, McKean–Vlasov equation.
1 Introduction
Variational inference (VI) is a popular Bayesian inference method for approximating intractable posterior distributions (Jordan et al., 1999; Bishop, 2006; Wainwright et al., 2008; Blei et al., 2017). The premise of VI is to perform such approximation via a constrained optimization problem of the form
$\min_{q\in\mathcal{Q}}D_{\mathrm{KL}}(q\|\pi)\,,$ (1.1)
where $\pi\in\mathcal{P}(\mathbb{R}^{n})$ is the posterior distribution, $\mathcal{Q}\subseteq\mathcal{P}(\mathbb{R}^{n})$ is some set of probability distributions, and $D_{\mathrm{KL}}(\cdot\|\cdot)$ is the Kullback–Leibler divergence. A common choice for $\mathcal{Q}$ is the set of product measures, i.e., $\mathcal{Q}=\big{\{}\prod_{i=1}^{n}q_{i},\;q_{i}\in\mathcal{P}(\mathbb{R})\big{\}}$, which is called the mean-field family. The corresponding variational problem for this choice of $\mathcal{Q}$ is called mean-field variational inference (MFVI). Compared to Markov chain Monte Carlo (MCMC) methods, which can have long burn-in periods, the optimization approach of VI allows for faster and more scalable inference in practice (Blei et al., 2017).
Optimization problems over spaces of probability distributions can in theory be solved with a gradient flow by endowing the space of probability distributions with the 2-Wasserstein metric $\mathcal{W}_{2}$ and using a discrete stepping scheme (Ambrosio et al., 2005; Ambrosio and Savaré, 2007; Santambrogio, 2017). The use of $\mathcal{W}_{2}$–gradient flows to solve the variational problem (1.1) has led to developments in theoretical understanding of VI (Lambert et al., 2022; Yao and Yang, 2022; Diao et al., 2023; Lacker, 2023) as well as the unconstrained version of (1.1) whose solution is the exact posterior (Trillos and Sanz-Alonso, 2020). Furthermore, a seminal result (Jordan et al., 1998) introduces a close connection between $\mathcal{W}_{2}$–gradient flows and PDEs: the discrete iterative scheme that describes the $\mathcal{W}_{2}$–gradient flow converges, in the small step-size limit, to the solution of a Fokker–Planck equation (Jordan et al., 1998, Theorem 5.1). The solution to the Fokker–Planck equation, in turn, describes the law of a particle evolving according to a stochastic differential equation (SDE). This PDE-and-SDE characterization of the $\mathcal{W}_{2}$–gradient flow opens up avenues for computation: it is generally difficult to compute Wasserstein distances, and thus gradient flows in the metric space, but the corresponding PDE and SDE are easier to approximate. Furthermore, the PDE and SDE are more amenable to theoretical analysis and convergence guarantees.
Since VI operates in a subset $\mathcal{Q}$ of the $\mathcal{W}_{2}$ space, efforts have been made to describe variants of VI—namely, VI with different choices of $\mathcal{Q}$—in terms of gradient flows. The difficulty lies in showing that a gradient flow constrained to a subspace of the $\mathcal{W}_{2}$ space is still a gradient flow. In other words, the challenge is to prove that the gradient flow is closed with respect to the subspace $\mathcal{Q}$. When this is shown, theoretical guarantees follow (Lambert et al., 2022; Yao and Yang, 2022; Lacker, 2023). In the case of MFVI, a recent paper (Lacker, 2023) shows that MFVI can be described by a $\mathcal{W}_{2}$–gradient flow constrained to the submanifold of product measures endowed with the $\mathcal{W}_{2}$ distance. The paper also shows that this gradient flow corresponds to a McKean–Vlasov equation obtained by an independent projection of the corresponding Langevin dynamics. The law of this McKean–Vlasov equation evolves according to a nonlinear Fokker–Planck equation, which is a system of coupled PDEs. This PDE-and-SDE description allows Lacker (2023) to specify convergence rates of the gradient flow to the solution of (1.1) as time goes to $\infty$.
Contributions
In this paper, we introduce a particle-based algorithm for MFVI, which we term PArticle VI (PAVI). The algorithm is built upon the SDE description of MFVI and stochastic approximation. Under standard regularity assumptions, we provide an upper bound on the 2-Wasserstein distance between the empirical measure of the particles at each iteration of Algorithm 2 and the solution of (1.1). Notably, our convergence guarantee is non-asymptotic and holds for any number of particles and total time duration. To our knowledge, this is the first end-to-end guarantee for particle-based MFVI.
The theoretical challenge lies not only in proving the convergence guarantees but also in designing appropriate time-discretization and particle approximation schemes for the McKean–Vlasov SDE. Other particle-based VI algorithms have been proposed in Yao and Yang (2022), but their theoretical results include error terms that do not diminish in the long-time limit and do not analyze the error stemming from the particle approximation. Proposing an implementable and efficient algorithm for the McKean–Vlasov SDE is not straightforward: as discussed in (Lacker, 2023, Section 2.3), the most obvious discretization does not contain a dissipative drift (which is needed for establishing uniform-in-time convergence rates).
Our algorithm not only has strong convergence guarantees (as discussed earlier) but is also nonparametric and flexible. The traditional algorithm for implementing MFVI, called coordinate ascent variational inference (CAVI), requires conditional conjugacy of the model parameters’ distributions (Bishop, 2006; Blei et al., 2017); in particular, the model’s complete conditionals must belong to exponential families. Our proposed method is both nonparametric and theoretically sound, like MCMC, while also being scalable and efficient thanks to the mean-field family of product measures, similarly to CAVI. On the mathematical side, we contribute novel results and techniques to the convergence analysis of McKean–Vlasov systems.
Related work
We outline recent theoretical advances in VI based on gradient flows in the $\mathcal{W}_{2}$ space as follows. When $\mathcal{Q}$ is the set of Gaussians, the resulting form of VI is called Gaussian VI. The set of Gaussians endowed with the $\mathcal{W}_{2}$ distance is also called the Bures–Wasserstein (BW) space (Bures, 1969), which is a subset of the $\mathcal{W}_{2}$ space. By studying $\mathcal{W}_{2}$–gradient flows constrained to the BW space, Lambert et al. (2022) propose a stochastic gradient descent algorithm for Gaussian VI—which discretizes the gradient flow—and provide theoretical guarantees, and Diao et al. (2023) develop a forward-backward algorithm along with convergence rates for Gaussian VI. These works suggest extensions to the case when $\mathcal{Q}$ is the set of mixtures of Gaussians, which corresponds to an essentially nonparametric $\mathcal{Q}$. See Yi and Liu (2023); Huix et al. (2024) for further developments.
In the mean-field case where $\mathcal{Q}$ consists of product distributions, Yao and Yang (2022) establish convergence guarantees for a coordinate-wise $\mathcal{W}_{2}$–gradient flow for MFVI in Bayesian models, and they propose two algorithms for practical implementation: a particle algorithm based on the Langevin SDE, and a function approximation method using neural networks. Their particle algorithm suffers from cumulative numerical errors that persist in the long-time limit (see (Yao and Yang, 2022, Lemma H.1 and Theorem D.2)), and they do not analyze the error stemming from the particle approximation. Lacker (2023) shows that MFVI can be described by a $\mathcal{W}_{2}$–gradient flow constrained to the submanifold of product distributions and establishes continuous-time convergence guarantees. Jiang et al. (2023) propose approximating the MFVI solution through optimization over polyhedral subsets of the Wasserstein space, and they introduce projected gradient descent algorithms with theoretical guarantees. Tran et al. (2023) also investigate particle MFVI; however, their theoretical analysis contains errors (see the footnote on (Jiang et al., 2023, Page 14)). Besides the above-mentioned algorithmic and computational developments for MFVI, there has also been recent work on the statistical accuracy of the solution to the MFVI problem for approximating the true posterior $\pi$. We refer the reader to, for example, Zhang and Gao (2020); Mukherjee and Sen (2022); Lacker et al. (2024) and the references therein for details.
Our convergence criteria resemble those from the literature on Langevin Monte Carlo (LMC), see Dalalyan (2017b); Durmus and Moulines (2017); Cheng and Bartlett (2018); Durmus et al. (2019); Durmus and Moulines (2019); Balasubramanian et al. (2022). LMC simulates an SDE whose stationary distribution is proportional to $e^{-V(\cdot)}$, where $V$ is a given potential. Non-asymptotic theoretical guarantees for LMC are given under the condition that the potential $V$ is continuously differentiable, strongly convex, and has Lipschitz gradient; see Durmus and Moulines (2019) for convergence guarantees in the $\mathcal{W}_{2}$ distance. We refer the reader to Chewi (2023) for a survey on LMC.
Our algorithm is built on the framework of gradient flows and particle algorithms. Gradient flows and particle algorithms have recently been applied to a wide range of problems in statistics and optimization. For example, Saeedi et al. (2017) propose a particle algorithm for approximating discrete distributions with the approximating family given by distributions with at most $K$ atoms (where $K$ is fixed). Cheng et al. (2024) propose particle algorithms for distribution approximation using generalized Wasserstein gradient flow of the Kullback-Leibler divergence. Their algorithms do not impose explicit specification of an approximating family $\mathcal{Q}$ and rely on neural networks to estimate certain vector fields used for updating. Their theoretical analysis relies on assumptions on neural network approximation accuracy and does not consider the error arising from particle approximation. We also refer the reader to Liu and Wang (2016); Gallouët and Monsaingeon (2017); Chizat and Bach (2018); Liu et al. (2019); Lu et al. (2019); Korba et al. (2020); Wang and Li (2020, 2022); Lu et al. (2023); Kook et al. (2024); Liu et al. (2024); Yan et al. (2024); Yao et al. (2024) and the references therein for further developments.
Outline
In Section 2, we state the problem setup, and discuss some relevant properties of MFVI. In Section 3, we lay out the McKean–Vlasov equation corresponding to the MFVI problem, and introduce a particle algorithm. In Section 4, we state our main theorem which provides a convergence guarantee for the stochastic particle algorithm, Algorithm 2. In Section 5, we prove the theoretical results stated in Section 4. Some technical lemmas used in the proof are given in Section 6. Section 7 is devoted to discussions.
Notation
We set up some notations that will be used throughout the paper. For any $d\in\mathbb{Z}_{+}=\{1,2,\cdots\}$, we denote $[d]=\{1,2,\cdots,d\}$ and let $\bm{I}_{d}$ represent the identity matrix of size $d$. For any $d\in\mathbb{Z}_{+}$, we denote by $\mathcal{P}(\mathbb{R}^{d})$ the set of probability distributions on $\mathbb{R}^{d}$, and by $\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{d})$ the family of probability distributions on $\mathbb{R}^{d}$ that are absolutely continuous with respect to the Lebesgue measure on $\mathbb{R}^{d}$. We denote the Euclidean norm by $\|\cdot\|_{2}$. For any two symmetric matrices $\bm{A}$ and $\bm{B}$ of the same size, we write $\bm{A}\preceq\bm{B}$ or $\bm{B}\succeq\bm{A}$ if $\bm{B}-\bm{A}$ is positive semidefinite. For any random variable $X$, we denote the law of $X$ by $\mathrm{Law}(X)$. For any $\bm{\mu}\in\mathbb{R}^{d}$ (where $d\in\mathbb{Z}_{+}$) and $d\times d$ positive definite matrix $\bm{\Sigma}$, we denote by $N(\bm{\mu},\bm{\Sigma})$ the multivariate normal distribution with mean vector $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$. For any two quantities $A,B>0$, we write $A=\mathcal{O}_{K}(B)$ if $A\leq CB$ for a positive constant $C$ that only depends on $K$. For any function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ (where $d\in\mathbb{Z}_{+}$), we denote its gradient and Hessian matrix by $\nabla f$ and $\nabla^{2}f$, respectively, and denote $\partial_{i}f(\bm{x}):=\frac{\partial f(\bm{x})}{\partial x_{i}}$ for any $\bm{x}=(x_{1},\cdots,x_{d})\in\mathbb{R}^{d}$ and $i\in[d]$.
2 Mean-field variational inference
In this section, we introduce the optimization problem involved in mean-field variational inference, and discuss relevant properties of this optimization problem.
2.1 Problem setup
We define the mean-field family of probability distributions as
$\mathcal{Q}=\bigg{\{}q^{1}\otimes\cdots\otimes q^{m}:~{}q^{i}\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}),~{}~{}\forall i\in[m]\bigg{\}}\,.$ (2.1)
We see that $\mathcal{Q}\subseteq\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{m})$ consists of product distributions. Given a distribution $p\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{m})$ whose density function is known up to a normalizing constant, mean-field variational inference (MFVI) aims to approximate $p$ by a product distribution in $\mathcal{Q}$, by solving the problem
$\min_{q\in\mathcal{Q}}D_{\mathrm{KL}}(q\|p)\,.$ (2.2)
Here $D_{\mathrm{KL}}(q\|p)=-\mathbb{E}_{q}\log(\frac{\mathrm{d}p}{\mathrm{d}q})$ is the Kullback–Leibler divergence (also known as the relative entropy) from $p$ to $q$. An optimal solution to (2.2) is called a mean-field approximation of $p$. MFVI is usually implemented using an algorithm called coordinate ascent variational inference (CAVI). In addition to the mean-field assumption, CAVI requires each complete conditional of the underlying model to belong to an exponential family (Bishop, 2006; Blei et al., 2017). To improve the expressive power of the variational family, nonparametric variants of CAVI have been proposed. For instance, $\mathcal{Q}$ can be the set of Gaussian mixtures (Gershman et al., 2012; Lambert et al., 2022). In this paper, we work with the entire mean-field family (2.1).
2.2 Key properties of MFVI
In the following, we discuss several key properties of the optimization problem (2.2). Throughout the rest of this paper, we assume that the target distribution $p\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{m})$ has density $Z^{-1}e^{-V(\cdot)}$, where $V:\mathbb{R}^{m}\rightarrow\mathbb{R}$ is a potential function and $Z$ is a normalizing constant.
We first introduce the following definitions.
Definition 2.1 (Entropy).
For any $d\in\mathbb{Z}_{+}$ and $\mu\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{d})$, we define the entropy of $\mu$ as
$H(\mu)=-\int_{\mathbb{R}^{d}}\mu(\bm{x})\log\mu(\bm{x})\mathrm{d}\bm{x}.$
Definition 2.2 (Marginals).
For any $\bm{x}=(x_{1},\cdots,x_{m})\in\mathbb{R}^{m}$ and $i\in[m]$, define
$\bm{x}_{-i}=(x_{1},\cdots,x_{i-1},x_{i+1},\cdots,x_{m}).$
For any $q=q^{1}\otimes\cdots q^{m}\in\mathcal{Q}$ and $i\in[m]$, define
$q^{-i}=q^{1}\otimes\cdots q^{i-1}\otimes q^{i+1}\otimes\cdots q^{m}.$
In other words, for a particle $\boldsymbol{x}$ or a distribution $q$, the subscript or superscript $-i$ means “all components except for the $i$-th one”.
Definition 2.3 (Approximate potential).
For any $\mu\in\mathcal{P}(\mathbb{R}^{m-1})$ and $i\in[m]$, let
$\bar{V}_{i}(\,\cdot\,,\mu)=\mathbb{E}_{\bm{x}_{-i}\sim\mu}V(x_{1},\cdots,x_{i-1},\,\cdot\,,x_{i+1},\cdots,x_{m}).$ (2.3)
The function $\bar{V}_{i}$ is an approximation (using a single component and a distribution $\mu$) to the potential function $V:\mathbb{R}^{m}\rightarrow\mathbb{R}$. It will be crucial for defining the mean-field particle dynamics. In this paper, any derivative of $\bar{V}_{i}$ is taken with respect to its first argument.
For any distribution $q\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{m})$, the Kullback–Leibler divergence from $p$ to $q$ can be rewritten as
$D_{\mathrm{KL}}(q\|p)=-\mathbb{E}_{\bm{x}\sim q}\log\bigg{(}\frac{\mathrm{d}p}{\mathrm{d}q}(\bm{x})\bigg{)}=\mathbb{E}_{\bm{x}\sim q}V(\bm{x})-H(q)+\log(Z).$
Note that for a product distribution $q\in\mathcal{Q}$, we have $\mathbb{E}_{\bm{x}\sim q}V(\bm{x})=\mathbb{E}_{x_{i}\sim q^{i}}\bar{V}_{i}(x_{i},q^{-i})$ and $H(q)=H(q^{i})+H(q^{-i})$. Therefore, for $q\in\mathcal{Q}$, the Kullback–Leibler divergence can be written as a function of the $i$-th component $q^{i}$ as
$D_{\mathrm{KL}}(q\|p)=\mathbb{E}_{x_{i}\sim q^{i}}\bar{V}_{i}(x_{i},q^{-i})-H(q^{i})-H(q^{-i})+\log(Z),\qquad\forall i\in[m]\,.$
For each $i\in[m]$, define the transform ${\mathcal{T}}_{i}:\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{m-1})\rightarrow\mathcal{P}_{\mathrm{AC}}(\mathbb{R})$ as
${\mathcal{T}}_{i}(\mu)=\frac{e^{-\bar{V}_{i}(\,\cdot\,,\mu)}}{\int_{\mathbb{R}}e^{-\bar{V}_{i}(x,\mu)}\mathrm{d}x}\,,$
which is an approximation of $p$ when all but one of the components of the argument of $V$ are distributed according to $\mu$. Using this notation, we have
$D_{\mathrm{KL}}(q\|p)=D_{\mathrm{KL}}\left(q^{i}\|{\mathcal{T}}_{i}(q^{-i})\right)-H(q^{-i})+\log(Z)-\log\bigg{(}\int_{\mathbb{R}}e^{-\bar{V}_{i}(x,q^{-i})}\mathrm{d}x\bigg{)}.$ (2.4)
If we minimize the Kullback–Leibler divergence $D_{\mathrm{KL}}(q\|p)$ while keeping $q^{-i}$ fixed, then the optimal $q^{i}$ is clearly ${\mathcal{T}}_{i}(q^{-i})$. Hence, the optimality of $q_{*}=\mathop{\mathrm{argmin}}_{q\in\mathcal{Q}}D_{\mathrm{KL}}(q\|p)$ implies the optimality of the $i$-th component $q_{*}^{i}={\mathcal{T}}_{i}(q_{*}^{-i})$, $\forall i\in[m]$. Thus, we get the following result, which characterizes the optimal $q_{*}$ as the fixed point of a transformation.
Lemma 2.1 (Fixed point characterization).
$q_{*}$ is the fixed point of transform ${\mathcal{T}}:~{}\mathcal{Q}\to\mathcal{Q}$, $q\mapsto{\mathcal{T}}_{1}(q^{-1})\otimes\cdots\otimes{\mathcal{T}}_{m}(q^{-m})$.
3 A particle algorithm for mean-field variational inference
In this section, we introduce a particle algorithm for mean-field variational inference, which we term PArticle VI (PAVI). In Section 3.1, we formulate the variational inference problem in terms of Fokker–Planck equations and the corresponding diffusion processes. Then, based on this formulation, we introduce the particle algorithm for practical implementation.
3.1 Fokker–Planck equations and diffusion processes
We start by considering a sub-problem of (2.2):
  • How to optimize $D_{\mathrm{KL}}(q\|p)$ with respect to $q^{i}$ while keeping $q^{-i}$ fixed?
In light of (2.4), this sub-problem amounts to solving for a univariate distribution ${\mathcal{T}}_{i}(q^{-i})\propto e^{-\bar{V}_{i}(\,\cdot\,,q^{-i})}$. We fix $q^{-i}$, and define
$f(\,\cdot\,)=\bar{V}_{i}(\,\cdot\,,q^{-i})\,.$
According to the general results by Jordan et al. (1998), the gradient flow for the Kullback–Leibler divergence $D_{\mathrm{KL}}(\,\cdot\,\|{\mathcal{T}}_{i}(q^{-i}))$ with respect to the 2-Wasserstein metric is the solution of the Fokker–Planck equation
$\begin{cases}&\partial_{t}\rho(x,t)=\partial_{x}[f^{\prime}(x)\rho(x,t)]+\partial^{2}_{xx}\rho(x,t),\\ &\rho(\,\cdot\,,0)=q^{i}(\,\cdot\,).\end{cases}$ (3.1)
Below is a formal definition of the aforementioned metric, which will be used throughout our analysis.
Definition 3.1 (2-Wasserstein metric).
For any $d\in\mathbb{Z}_{+}$ and $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$, the 2-Wasserstein distance between $\mu$ and $\nu$ is
$\mathcal{W}_{2}(\mu,\nu)=\inf_{\gamma\in\mathcal{C}(\mu,\nu)}\sqrt{\mathbb{E}_{(\bm{x},\bm{y})\sim\gamma}\|\bm{x}-\bm{y}\|_{2}^{2}}\,,$
where $\mathcal{C}(\mu,\nu)$ is the set of joint distributions on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with $\mu$ and $\nu$ as their marginals. Any $\gamma\in\mathcal{C}(\mu,\nu)$ is called a coupling between $\mu$ and $\nu$.
The Fokker–Planck equation is deeply connected to a stochastic differential equation: the solution $\rho(\,\cdot\,,t)$ of (3.1) is the law of the diffusion process
$\begin{cases}&\mathrm{d}Y_{t}=-f^{\prime}(Y_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},\\ &Y_{0}\sim q^{i},\end{cases}$ (3.2)
where $Y_{t}:[0,\infty)\rightarrow\mathbb{R}$ and $\{B_{t}\}_{t\geq 0}$ is a standard Brownian motion that is independent of $Y_{0}$.
Simultaneous updates of all the marginal distributions $\{q^{i}\}_{i=1}^{m}$, where $q^{-i}$ is fixed for each update, lead to a McKean–Vlasov process:
$\begin{cases}&\mathrm{d}Y_{t}^{i}=-\bar{V}_{i}^{\prime}(Y_{t}^{i},\mathrm{Law}(\bm{Y}_{t}^{-i}))\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}^{i},\qquad i\in[m],\,t\geq 0\\ &\bm{Y}_{0}\sim q^{1}\otimes\cdots\otimes q^{m}\,,\end{cases}.$ (3.3)
where $\bar{V}_{i}$ is defined as in (2.3), $\bm{Y}_{t}=(Y_{t}^{1},\cdots,Y_{t}^{m})$, and $\{(B_{t}^{1},\cdots,B_{t}^{m})\}_{t\geq 0}$ is an $m$-dimensional standard Brownian motion that is independent of $\bm{Y}_{0}$. The trajectory $t\mapsto\mathrm{Law}(\bm{Y}_{t})$ is the Wasserstein gradient flow for optimizing $D_{\mathrm{KL}}(\,\cdot\,\|p)$ over the mean-field family $\mathcal{Q}$ (Lacker, 2023). Note that the process $\{\bm{Y}_{t}\}_{t\geq 0}$ has independent coordinates because the $m$-dimensional Brownian motion has independent coordinates and the other coordinates $\bm{Y}_{t}^{-i}$ are integrated out in the drift term.
3.2 PArticle VI
We are ready to present PArticle VI (PAVI), a particle algorithm for mean-field variational inference. To set the stage, we first introduce some convenient notations.
Definition 3.2 (Empirical measure of an array of particles).
Let $S_{N}$ be the set of all permutations of $[N]$. For any array $\bm{A}\in\mathbb{R}^{m\times N}$, denote by $A^{i,j}$ its $(i,j)$-th entry, and let $\bm{A}^{i,:}=(A^{i,1},\cdots,A^{i,N})$ be the $i$-th row, $\bm{A}^{:,j}=(A^{1,j},\cdots,A^{m,j})$ be the $j$-th column, and $\bm{A}^{-i,j}=(A^{1,j},\cdots,A^{i-1,j},A^{i+1,j},\cdots,A^{m,j})$ be the $j$-th column of $\bm{A}$ with the $i$-th entry removed. For each $i\in[m]$, let $q_{\bm{A}}^{i}=\frac{1}{N}\sum_{j=1}^{N}\delta_{A^{i,j}}$. Define $q_{\bm{A}}=q_{\bm{A}}^{1}\otimes\cdots\otimes q_{\bm{A}}^{m}$.
Remark 1.
For each $i\in[m]$, the measure $q^{i}_{\bm{A}}$ is the empirical measure of $\bm{A}^{i,:}$, the $i$-th component for the $N$ particles. $q_{\bm{A}}\in\mathcal{P}(\mathbb{R}^{m})$ is the product measure with $\{q^{i}_{\bm{A}}\}_{i=1}^{m}$ as its marginals. Hence, it has $N^{m}$ atoms in total.
Our first version of PAVI, Algorithm 1, can be viewed as a direct discretization of the McKean–Vlasov process (3.3).
Input: Function $V:~{}\mathbb{R}^{m}\to\mathbb{R}$, initial value $\bm{X}_{0}\in\mathbb{R}^{m\times N}$, step-size $h>0$, number of iterations $T\in\mathbb{Z}_{+}$.
For $n=0,1,\ldots,T-1$:
       For $i=1,\cdots,m$:
            Draw a random vector $\bm{\xi}_{n}^{i,:}$ from $N(0,\bm{I}_{N})$, the $N$-dimensional standard normal distribution.
            For $j=1,\cdots,N$:
                 Compute $X_{n+1}^{i,j}=X_{n}^{i,j}-h\bar{V}_{i}^{\prime}(X_{n}^{i,j},q^{-i}_{\bm{X}_{n}})+\sqrt{2h}\xi_{n}^{i,j}$.
Output: $\bm{X}_{T}$.
Algorithm 1 A gradient descent version of PAVI
If the columns of $\bm{X}_{0}$ are i.i.d. samples from a probability distribution over $\mathbb{R}^{m}$, $N\to\infty$ and $h\to 0$, then Algorithm 1 becomes the Wasserstein gradient flow in Lacker (2023). Meanwhile, we note that Algorithm 1 is intuitive but impractical: the gradient evaluation is prohibitively expensive due to the cost of averaging over $N^{m-1}$ terms. For computational reasons, we propose a second version of PAVI, presented in Algorithm 2 below, which involves a stochastic approximation of the gradient term.
Input: Function $V:~{}\mathbb{R}^{m}\to\mathbb{R}$, initial value $\bm{X}_{0}\in\mathbb{R}^{m\times N}$, step-size $h>0$, batch size $B\in\mathbb{Z}_{+}$, number of iterations $T\in\mathbb{Z}_{+}$.
For $n=0,1,\ldots,T-1$:
       Draw $B$ i.i.d. samples $\{\bm{z}_{n}^{:,b}\}_{b=1}^{B}$ from $q_{\bm{X}_{n}}$.
       For $i=1,\cdots,m$:   # can be parallelized
            Define a function $g_{n}^{i}(\,\cdot\,)=\frac{1}{B}\sum_{b=1}^{B}\partial_{i}V(z_{n}^{1,b},\cdots,z_{n}^{(i-1),b},\cdot,z_{n}^{(i+1),b},\cdots,z_{n}^{m,b})$.
            Draw a random vector $\bm{\xi}_{n}^{i,:}$ from $N(0,\bm{I}_{N})$, the $N$-dimensional standard normal distribution.
            For $j=1,\cdots,N$:
                 Compute $X_{n+1}^{i,j}=X_{n}^{i,j}-hg_{n}^{i}(X_{n}^{i,j})+\sqrt{2h}\xi_{n}^{i,j}$.
Output: $\bm{X}_{T}$.
Algorithm 2 PAVI
Algorithm 1 is the limiting case of Algorithm 2 (PAVI) as $B\to\infty$. The finite batch size in PAVI ensures its computational efficiency. The output $\bm{X}_{T}$ from Algorithm 2 is associated with its empirical distribution $q_{\bm{X}_{T}}$, which is a product distribution that approximates $q_{*}$, the solution to (2.2). In the next section, we will prove the convergence of $q_{\bm{X}_{T}}$ towards $q_{*}$ in the $\mathcal{W}_{2}$ distance. Our theory will also provide practical guidance on choosing $B$. It is worth pointing out that convergence guarantees for Algorithm 1 can be obtained by taking the limit $B\to\infty$.
4 Theoretical analysis
In this section, we provide a theoretical analysis of Algorithm 2, the particle algorithm with stochastic approximation introduced in Section 3. Recall from Section 2.2 that the target distribution $p\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{m})$ has density $Z^{-1}e^{-V(\cdot)}$. We work under the following two regularity assumptions.
Assumption 4.1 (Strong convexity and smoothness).
There exist constants $0<\alpha\leq L<\infty$ such that for any $\bm{x}\in\mathbb{R}^{m}$, we have $\alpha\bm{I}_{m}\preceq\nabla^{2}V(\bm{x})\preceq L\bm{I}_{m}$.
Assumption 4.2 (Bounded third-order derivatives).
For any $\bm{x}\in\mathbb{R}^{m}$ and $i\in[m]$, we have $\left|\frac{\partial^{3}}{\partial x_{i}^{3}}V(\bm{x})\right|\leq\widetilde{L}<\infty$.
Under Assumption 4.1, we have the following result from Lacker (2023); Lacker et al. (2024), which shows the existence and uniqueness of the MFVI solution $q_{*}$ and establishes the exponential convergence of the law of $\bm{Y}_{t}$ (recall (3.3)) towards $q_{*}$.
Lemma 4.1 (Theorems 1.1 of Lacker et al. (2024) and Theorem 2.3 of Lacker (2023)).
Under Assumption 4.1, the following holds:
  • (a)
    The optimization problem (2.2) has a unique solution $q_{*}$.
  • (b)
    For any $t\geq 0$, we have
    $\mathcal{W}_{2}(\mathrm{Law}(\bm{Y}_{t}),q_{*})\leq e^{-\alpha t}\mathcal{W}_{2}(\mathrm{Law}(\bm{Y}_{0}),q_{*}).$
Lemma 4.1 focuses on the long-time convergence of the continuous-time process $\bm{Y}_{t}$ but does not provide algorithmic guarantees. Our main theoretical result, presented in Theorem 4.1 below, establishes the first end-to-end theoretical guarantee for an implementable particle algorithm (Algorithm 2) for MFVI. In Theorem 4.1, we provide a non-asymptotic upper bound on the $\mathcal{W}_{2}$ distance between $q_{\bm{X}_{n}}$ (the empirical distribution of the $N$ particles at the $n$-th iteration) and the MFVI solution $q_{*}$.
Theorem 4.1.
Suppose that Assumptions 4.1 and 4.2 hold. Consider Algorithm 2 with $N\geq 2$ and $0<h<\min\{2/(\alpha+L),B\alpha/(4L^{2})\}$. For any $T\in\mathbb{Z}_{+}$ and $10\leq n\leq T-$ 1, we have
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq(1-\alpha h/4)^{n}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}$
$\quad+\frac{C\sqrt{m}}{\alpha}\bigg{(}h^{-1}\sqrt{\frac{\log N}{\alpha N}}+\sqrt{h\max\{h,(LB)^{-1}\}(L^{4}/\alpha+\widetilde{L}^{2})}\bigg{)},$
where $C>0$ is a universal constant. In particular, when $B\geq 1/(Lh)$, we have
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq(1-\alpha h/4)^{n}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}+\frac{C\sqrt{m}}{\alpha^{3/2}}\bigg{(}h^{-1}\sqrt{\frac{\log N}{N}}+h\sqrt{L^{4}+\widetilde{L}^{2}\alpha}\bigg{)}.$ (4.1)
Recall that as $B\to\infty$, Algorithm 2 reduces to Algorithm 1. Therefore, the error bound (4.1) also holds for Algorithm 1. The proof of Theorem 4.1 will be presented in Section 5.
By taking $h=1/(LN^{1/4})$ and $B=\lceil 1/(Lh)\rceil$ in Theorem 4.1, we obtain the following corollary.
Corollary 4.1.
Let $h=1/(LN^{1/4})$ and $B=\lceil 1/(Lh)\rceil$. We have
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq(1-\alpha h/4)^{n}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}+\frac{\sqrt{m}}{N^{1/4}}\cdot\frac{CL}{\alpha^{3/2}}\bigg{(}\sqrt{\log N}+\sqrt{1+\widetilde{L}^{2}\alpha/L^{4}}\bigg{)},$
where $C>0$ is a universal constant.
We note that the non-asymptotic upper bound in Corollary 4.1 holds for any number of particles $N$ and any number of iterations $n$, and it converges to 00 as $N$ and $n$ grow at appropriate rates (see below for a detailed analysis of the required sizes of $N$ and $n$ to achieve a given precision $\varepsilon\in(0,1)$). Thus, our result establishes the long-time convergence of Algorithm 2. To the best of our knowledge, this is the first such convergence guarantee for particle-based MFVI. The particle algorithm proposed in Yao and Yang (2022) has cumulative numerical errors that do not diminish even when the number of particles and the number of iterations grow to infinity. Furthermore, the theoretical analysis therein does not address errors arising from the particle approximation.
By Corollary 4.1, if we take the step-size $h=1/(LN^{1/4})$ and the batch size $B=\lceil 1/(Lh)\rceil$, then for any $\varepsilon\in(0,1)$, as long as the number of particles $N$ satisfies $N\geq Km^{2}(\log(m/\varepsilon^{2}))^{2}\varepsilon^{-4}$ (where $K$ is a sufficiently large constant depending only on $\alpha,L,\tilde{L}$), the number of iterations $T$ required in order for $\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{T}},q_{*})}\leq\varepsilon$ is
$\mathcal{O}_{\alpha,L}\bigg{(}N^{1/4}\max\Big{\{}\log\Big{(}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}\varepsilon^{-1}\Big{)},1\Big{\}}\bigg{)}.$
In particular, if the number of particles $N$ is taken to be of order $m^{2}(\log(m/\varepsilon^{2}))^{2}\varepsilon^{-4}$, the required number of iterations is
$\mathcal{O}_{\alpha,L,\tilde{L}}\bigg{(}\sqrt{m\log(m/\epsilon^{2})}\varepsilon^{-1}\max\Big{\{}\log\Big{(}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}\varepsilon^{-1}\Big{)},1\Big{\}}\bigg{)}.$ (4.2)
Notably, the requirement (4.2) on the number of iterations grows mildly with the dimension $m$ and does not suffer from the curse of dimensionality.
We also note that our algorithm, Algorithm 2, provides a direct approximation $q_{\bm{X}_{n}}$ to the MFVI solution $q_{*}$, with its approximation accuracy guaranteed by Theorem 4.1 and Corollary 4.1. Langevin Monte Carlo (LMC), by contrast, tracks the evolution of a single particle. Theoretical results for LMC (Dalalyan, 2017b; Durmus and Moulines, 2017; Cheng and Bartlett, 2018; Durmus et al., 2019; Durmus and Moulines, 2019) primarily focus on bounding the $\mathcal{W}_{2}$ distance (or other metrics on probability measures) between the distribution of this particle at the $n$-th iteration and the target distribution. In order to approximate the target distribution using the output of LMC, one needs to generate many such particles and use their empirical distribution. One option is to run $T_{0}$ independent replications of LMC, each time collecting the particle $\mathbf{x}_{t,n}\in\mathbb{R}^{m}$ (where $t\in[T_{0}]$) at the $n$-th iteration, and use the empirical measure $T_{0}^{-1}\sum_{t=1}^{T_{0}}\delta_{\mathbf{x}_{t,n}}$ to approximate the target distribution. Note that this approach would suffer from the curse of dimensionality: in high dimensions, the $\mathcal{W}_{2}$ distance between the empirical distribution and the true one is known to converge very slowly as the sample size $T_{0}$ increases (Weed and Bach, 2019).
5 Proof of Theorem 4.1
In this section, we present the proof of Theorem 4.1.
5.1 Preparations
The main idea is to couple the iterates of Algorithm 2 with a stationary process. First, we introduce a metric for measuring the difference between two particle swarms.
Definition 5.1.
For two Borel probability distributions $\mu$ and $\nu$ on $\mathbb{R}^{m\times N}$, define $\mathcal{C}(\mu,\nu)$ as the set of all couplings between $\mu$ and $\nu$. Let
$\widetilde{\mathcal{W}}(\mu,\nu)=\inf_{\gamma\in\mathcal{C}(\mu,\nu)}\sqrt{\mathbb{E}_{(\bm{X},\bm{Y})\sim\gamma}\mathcal{W}_{2}^{2}(q_{\bm{X}},q_{\bm{Y}})}\,.$
Let $\bm{Z}$ and $\bm{W}$ be two random arrays in $\mathbb{R}^{m\times N}$, which induce random product distributions $q_{\bm{Z}}$ and $q_{\bm{W}}$ as in Definition 3.2, respectively. Intuitively, if $q_{\bm{W}}$ concentrates around a fixed probability distribution $\mu$, then $\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},q_{\bm{W}})$ should be close to $\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},\mu)$ and thus the former distance is insensitive to the dependence between $\bm{Z}$ and $\bm{W}$. In that case, $\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},q_{\bm{W}})}$ is close to the infimum of $\widetilde{\mathcal{W}}\left(\mathrm{Law}(\bm{Z}),\mathrm{Law}(\bm{W})\right)$ over all couplings between $\mathrm{Law}(\bm{Z})$ and $\mathrm{Law}(\bm{W})$. This is confirmed by the lemma below.
Lemma 5.1.
Let $\bm{Z},\bm{W}\in\mathbb{R}^{m\times N}$ be random. For any probability measure $\mu\in\mathcal{P}(\mathbb{R}^{m})$,
$\Big{|}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},q_{\bm{W}})}-\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},\mu)}\Big{|}\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{W}},\mu)}\,,$
$0\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},q_{\bm{W}})}-\widetilde{\mathcal{W}}(\mathrm{Law}(\bm{Z}),\mathrm{Law}(\bm{W}))\leq 2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{W}},\mu)}\,.$
Proof of Lemma 5.1.
By the triangle inequality,
$|\mathcal{W}_{2}(q_{\bm{Z}},q_{\bm{W}})-\mathcal{W}_{2}(q_{\bm{Z}},\mu)|\leq\mathcal{W}_{2}(q_{\bm{W}},\mu)\,.$
By Minkowski’s inequality,
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{W}},\mu)}\geq\sqrt{\mathbb{E}|\mathcal{W}_{2}(q_{\bm{Z}},q_{\bm{W}})-\mathcal{W}_{2}(q_{\bm{Z}},\mu)|^{2}}\geq\Big{|}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},q_{\bm{W}})}-\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},\mu)}\Big{|}\,.$
Therefore, for any coupling $\gamma$ between $\mathrm{Law}(\bm{Z})$ and $\mathrm{Law}(\bm{W}))$, the quantity $\sqrt{\mathbb{E}_{(\bm{Z}^{\prime},\bm{W}^{\prime})\sim\gamma}\mathcal{W}_{2}^{2}(q_{\bm{Z}^{\prime}},q_{\bm{W}^{\prime}})}$ belongs to a fixed interval of length $2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{W}},\mu)}$. This immediately proves the desired result. ∎
5.2 Main proof
Let $\bm{Y}\in\mathbb{R}^{m\times N}$ be a random array such that the $N$ particles $\{\bm{Y}^{:,j}\}_{j=1}^{N}$ are i.i.d. samples from $q_{*}$. When $N$ is large, $q_{\bm{Y}}$ concentrates around $q_{*}$ (Corollary 6.1).
In addition, let $\{B_{t}^{i,j}\}_{t\geq 0}$ for $(i,j)\in[m]\times[N]$ be i.i.d. standard Brownian motions. Suppose that $\bm{Y}$ and $\{\bm{B}_{t}\}_{t\geq 0}$ are independent, and both are independent of the randomness in Algorithm 2. For $(i,j)\in[m]\times[N]$, define $\{Y_{t}^{i,j}\}_{t\geq 0}$ as the unique strong solution to the stochastic differential equation
$\begin{cases}&\mathrm{d}Y_{t}^{i,j}=-\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}^{i,j}\\ &Y_{0}^{i,j}=Y^{i,j}\,.\end{cases}$
We present a useful lemma on the regularity of $\{\bar{V}_{i}\}_{i=1}^{m}$, which guarantees the well-posedness of the above SDE (Karatzas and Shreve, 1991). The proof is immediate and thus omitted.
Lemma 5.2.
Let Assumption 4.1 and 4.2 hold. For any $\mu\in\mathcal{P}(\mathbb{R}^{m-1})$ and $x\in\mathbb{R}$, we have $\alpha\leq\bar{V}_{i}^{\prime\prime}(x,\mu)\leq L$ and $|V_{i}^{\prime\prime\prime}(x,\mu)|\leq\widetilde{L}$.
By definition, we have
$Y_{h}^{i,j}=Y_{0}^{i,j}-\int_{0}^{h}\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})\mathrm{d}t+\sqrt{2}B_{h}^{i,j}\approx Y_{0}^{i,j}-h\bar{V}_{i}^{\prime}(Y_{0}^{i,j},q_{*}^{-i})\mathrm{d}t+\sqrt{2h}(B_{h}^{i,j}/\sqrt{h})$
and $B_{h}^{i,j}/\sqrt{h}\sim N(0,1)$. Hence, the update from $\bm{Y}_{0}$ to $\bm{Y}_{h}$ is similar to that from $\bm{X}_{n}$ to $\bm{X}_{n+1}$. This observation allows us to bound $\widetilde{\mathcal{W}}(\mathrm{Law}(\bm{X}_{n+1}),\mathrm{Law}(\bm{Y}_{h}))$ using $\widetilde{\mathcal{W}}(\mathrm{Law}(\bm{X}_{n}),\mathrm{Law}(\bm{Y}_{0}))$. For any $t\geq 0$, $\bm{Y}_{t}$ has the same distribution as $\bm{Y}$. We can establish the following recursive inequality. See Section 5.3 for its proof.
Lemma 5.3.
Let Assumptions 4.1 and 4.2 hold. Suppose that $0<h<\min\{2/(\alpha+L),B\alpha/(4L^{2})\}$. Define $W_{n}=\widetilde{\mathcal{W}}(\mathrm{Law}(\bm{X}_{n}),\mathrm{Law}(\bm{Y}))$ for $n\in\mathbb{N}$. We have
$W_{n+1}^{2}\leq\bigg{(}(1-\alpha h/4)W_{n}+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}\bigg{)}^{2}+4mh^{2}\max\{h,(LB)^{-1}\}(L^{4}/\alpha^{2}+\widetilde{L}^{2}/\alpha).$
Next, we invoke a useful lemma.
Lemma 5.4 (Lemma 1 in Dalalyan and Karagulyan (2019)).
Let $A$, $B$ and $C$ be non-negative numbers such that $A\in(0,1)$. Assume that a sequence of non-negative numbers $\{x_{n}\}_{n=0}^{\infty}$ satisfies the recursive inequality
$x_{n+1}^{2}\leq[(1-A)x_{n}+C]^{2}+B^{2},\qquad\forall n\in\mathbb{Z}_{+}.$
Then,
$x_{n}\leq(1-A)^{n}x_{0}+\frac{C}{A}+\frac{B}{\sqrt{A}},\qquad\forall n\in\mathbb{Z}_{+}.$
Applying Lemma 5.4 to the sequence $\{W_{n}\}_{n=0}^{\infty}$ in Lemma 5.3, we get
$W_{n}$ $\leq(1-\alpha h/4)^{n}W_{0}+\frac{2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}}{\alpha h/4}+\sqrt{\frac{4mh^{2}\max\{h,(LB)^{-1}\}(L^{4}/\alpha^{2}+\widetilde{L}^{2}/\alpha)}{\alpha h/4}}$
$\leq(1-\alpha h/4)^{n}W_{0}+\frac{8\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}}{\alpha h}+4\sqrt{mh\max\{h,(LB)^{-1}\}(L^{4}/\alpha^{3}+\widetilde{L}^{2}/\alpha^{2})}.$
By Lemma 5.1,
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})}+\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}$
$\leq\widetilde{\mathcal{W}}(\mathrm{Law}(\bm{X}_{n}),\mathrm{Law}(\bm{Y}))+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}+\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}$
$=W_{n}+3\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}.$
On the other hand,
$W_{0}=\widetilde{\mathcal{W}}(\mathrm{Law}(\bm{X}_{0}),\mathrm{Law}(\bm{Y}))\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{\bm{Y}})}\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}+\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}.$
Based on the above estimates, the assumption $h<2/(\alpha+L)$, and Corollary 6.1, there exists a universal constant $C$ such that
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq(1-\alpha h/4)^{n}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}$
$\quad+\frac{C\sqrt{m}}{\alpha}\bigg{(}h^{-1}\sqrt{\frac{\log N}{\alpha N}}+\sqrt{h\max\{h,(LB)^{-1}\}(L^{4}/\alpha+\widetilde{L}^{2})}\bigg{)}.$
It remains to prove Lemma 5.3.
5.3 Proof of Lemma 5.3
5.3.1 Step 1: Analysis of $W_{n}$
According to Lemma 5.1,
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})}\leq W_{n}+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}.$ (5.1)
We now look into the quantity $\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})$. Let $X_{n}^{i,(1)}\leq\cdots\leq X_{n}^{i,(N)}$ and $Y^{i,(1)}\leq\cdots\leq Y^{i,(N)}$ be the order statistics of $\{X_{n}^{i,j}\}_{j=1}^{N}$ and $\{Y^{i,j}\}_{j=1}^{N}$, respectively. Then,
$\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}}^{i},q_{\bm{Y}}^{i})=\frac{1}{N}\sum_{j=1}^{N}|X_{n}^{i,(j)}-Y^{i,(j)}|^{2},\qquad\forall i\in[m].$
Denote by $S_{N}$ the set of all permutations of $[N]$. There exist $\{\tau_{i}\}_{i=1}^{m}\subseteq S_{N}$, where $\tau_{i}$ is determined by $\bm{X}_{n}^{i,:}$ and $\bm{Y}^{i,:}$, such that
$\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}}^{i},q_{\bm{Y}}^{i})=\frac{1}{N}\sum_{j=1}^{N}|X_{n}^{i,\tau_{i}(j)}-Y^{i,j}|^{2},\qquad\forall i\in[m].$
For each $i\in[m]$, $\{Y^{i,j}\}_{j=1}^{N}$ are i.i.d. from an absolutely continuous distribution $q_{*}^{i}$. The continuity of $q_{*}^{i}$ and the symmetry of $\{Y^{i,j}\}_{j=1}^{N}$ imply the following.
Fact 5.1.
Conditioned on $\bm{X}_{n}$, the permutations $\{\tau_{i}\}_{i=1}^{m}$ are
  • uniquely defined with probability 1;
  • independent and uniformly distributed on $S_{N}$.
Define an array $\bm{U}$ with $U^{i,j}=X_{n}^{i,\tau_{i}(j)}$. Fact 5.1 implies that given $\bm{X}_{n}$, $\{\bm{U}^{:,j}\}_{j=1}^{N}$ are i.i.d. samples from $q_{\bm{X}_{n}}$. In addition, we have $q_{\bm{U}}=q_{\bm{X}_{n}}$ and
$\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}}^{i},q_{\bm{Y}}^{i})=\frac{1}{N}\sum_{j=1}^{N}|U^{i,j}-Y^{i,j}|^{2},\qquad\forall i\in[m].$
By Lemma 6.1,
$\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}}^{i},q_{\bm{Y}}^{i})=\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}|U^{i,j}-Y^{i,j}|^{2}.$ (5.2)
5.3.2 Step 2: Analysis of $W_{n+1}$
Define two arrays $\bm{M},\bm{W}\in\mathbb{R}^{m\times N}$ through
$M^{i,j}=U^{i,j}-hg_{i}(U^{i,j})+\sqrt{2}B_{h}^{i,j},$
$W^{i,j}=U^{i,j}-h\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{X}_{n}}^{-i})+\sqrt{2}B_{h}^{i,j}.$
It is easily seen that $q_{\bm{M}}$ and $q_{\bm{X}_{n+1}}$ have the same distribution, and $\bm{W}=\mathbb{E}(\bm{M}|\bm{X}_{n},\{\bm{Y}_{t}\}_{t=0}^{h},\{\bm{B}_{t}\}_{t=0}^{h})$.
For any $t\geq 0$, $\bm{Y}_{t}$ has the same distribution as $\bm{Y}$. Hence,
$W_{n+1}^{2}=\widetilde{\mathcal{W}}^{2}(\mathrm{Law}(\bm{X}_{n+1}),\mathrm{Law}(\bm{Y}))\leq\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{M}},q_{\bm{Y}_{h}})\leq\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|M^{i,j}-Y_{h}^{i,j}|^{2}.$
By $\bm{W}=\mathbb{E}(\bm{M}|\bm{X}_{n},\{\bm{Y}_{t}\}_{t=0}^{h},\{\bm{B}_{t}\}_{t=0}^{h})$, we have
$\mathbb{E}|M^{i,j}-Y_{h}^{i,j}|^{2}=\mathbb{E}[(M^{i,j}-W^{i,j})+(W^{i,j}-Y_{h}^{i,j})]^{2}=\mathbb{E}|M^{i,j}-W^{i,j}|^{2}+\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}.$
Hence,
$W_{n+1}^{2}\leq\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\Big{(}\mathbb{E}|M^{i,j}-W^{i,j}|^{2}+\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}\Big{)}.$ (5.3)
Below we look into the quantities on the right-hand side.
5.3.3 Step 3: Analysis of $\mathbb{E}|M^{i,j}-W^{i,j}|^{2}$
By definition,
$\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}|M^{i,j}-W^{i,j}|^{2}$ $=\frac{h^{2}}{N}\sum_{j=1}^{N}\mathbb{E}|g_{i}(U^{i,j})-\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{X}_{n}}^{-i})|^{2}$
$=\frac{h^{2}}{N}\sum_{j=1}^{N}\mathbb{E}|g_{i}(X_{n}^{i,\tau_{i}(j)})-\bar{V}_{i}^{\prime}(X_{n}^{i,\tau_{i}(j)},q_{\bm{X}_{n}}^{-i})|^{2}$
$=\frac{h^{2}}{N}\sum_{j=1}^{N}\mathbb{E}|g_{i}(X_{n}^{i,j})-\bar{V}_{i}^{\prime}(X_{n}^{i,j},q_{\bm{X}_{n}}^{-i})|^{2}.$
Let us focus on $i=1$ for a moment. Conditioned on $\bm{X}_{n}$, $\{\bm{z}_{n}^{-1,b}\}_{b=1}^{B}$ are i.i.d. samples from $q_{\bm{X}_{n}}^{-1}$ and
$g_{1}(X_{n}^{1,j})=\frac{1}{B}\sum_{b=1}^{B}\partial_{1}V(X_{n}^{1,j},\bm{z}_{n}^{-1,b}).$
Hence,
$\mathbb{E}\Big{(}|g_{1}(X_{n}^{1,j})-\bar{V}_{1}^{\prime}(X_{n}^{1,j},q_{\bm{X}_{n}}^{-1})|^{2}\Big{|}\bm{X}_{n}\Big{)}$ $=\operatorname{{\rm var}}[g_{1}(X_{n}^{1,j})|\bm{X}_{n}]=\frac{1}{B}\operatorname{{\rm var}}[\partial_{1}V(X_{n}^{1,j},\bm{z}_{n}^{-1,1})|\bm{X}_{n}]$
$\leq\frac{1}{B}\mathbb{E}\Big{\{}[\partial_{1}V(X_{n}^{1,j},\bm{z}_{n}^{-1,1})]^{2}\Big{|}\bm{X}_{n}\Big{\}}.$
Based on the above and the fact that $\bm{z}_{n}^{:,1}\sim q_{\bm{X}_{n}}$,
$\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}|M^{1,j}-W^{1,j}|^{2}\leq\frac{h^{2}}{BN}\sum_{j=1}^{N}\mathbb{E}[\partial_{1}V(X_{n}^{1,j},\bm{z}_{n}^{-1,1})]^{2}=\frac{h^{2}}{B}\mathbb{E}|\partial_{1}V(\bm{z}_{n}^{:,1})|^{2}.$
As a result,
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|M^{i,j}-W^{i,j}|^{2}\leq\frac{h^{2}}{B}\mathbb{E}\|\nabla V(\bm{z}_{n}^{:,1})\|_{2}^{2}=\frac{h^{2}}{B}\mathbb{E}_{\bm{X}_{n}}[\mathbb{E}_{\bm{z}\sim q_{\bm{X}_{n}}}\|\nabla V(\bm{z})\|_{2}^{2}].$ (5.4)
We now relate the right-hand side of (5.4) to $\mathbb{E}_{\bm{y}\sim q_{\bm{Y}}}\|\nabla V(\bm{y})\|_{2}^{2}$. We condition on $(\bm{X}_{n},\bm{Y})$ and denote by $\gamma$ an $\mathcal{W}_{2}$-optimal coupling between $q_{\bm{X}_{n}}$ and $q_{\bm{Y}}$ such that
$\mathbb{E}_{(\bm{u},\bm{v})\sim\gamma}\|\bm{u}-\bm{v}\|_{2}^{2}=\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}}).$
Construct random vectors $\bm{z},\bm{y}\in\mathbb{R}^{m}$ such that the conditional distribution of $(\bm{z},\bm{y})$ given $(\bm{X}_{n},\bm{Y})$ is $\gamma$. Then,
$\mathbb{E}\|\bm{z}-\bm{y}\|_{2}^{2}=\mathbb{E}[\mathbb{E}(\|\bm{z}-\bm{y}\|_{2}^{2}|\bm{X}_{n},\bm{Y})]=\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}}).$
By Assumption 4.1,
$\mathbb{E}\|\nabla V(\bm{z})\|_{2}^{2}$ $=\mathbb{E}\|[\nabla V(\bm{z})-\nabla V(\bm{y})]+\nabla V(\bm{y})\|_{2}^{2}$
$\leq 2\mathbb{E}\|\nabla V(\bm{z})-\nabla V(\bm{y})\|_{2}^{2}+2\mathbb{E}\|\nabla V(\bm{y})\|_{2}^{2}$
$\leq 2L^{2}\mathbb{E}\|\bm{z}-\bm{y}\|_{2}^{2}+2\mathbb{E}\|\nabla V(\bm{y})\|_{2}^{2}$
$\leq 2L^{2}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+2\mathbb{E}\|\nabla V(\bm{y})\|_{2}^{2}.$
The conditional distribution of $\bm{y}$ given $\bm{Y}$ is $q_{\bm{Y}}$. Hence, $\bm{y}\sim q_{*}$ and $\mathbb{E}\|\nabla V(\bm{y})\|_{2}^{2}=\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}\leq mL^{2}/\alpha$. Plugging these estimates into (5.4), we get
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|M^{i,j}-W^{i,j}|^{2}\leq\frac{2h^{2}L^{2}}{B}[\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+m/\alpha].$ (5.5)
5.3.4 Step 4: Analysis of $\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}$
Note that
$Y_{h}^{i,j}$ $=Y_{0}^{i,j}-\int_{0}^{h}\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})\mathrm{d}t+\sqrt{2}B_{h}^{i,j}$
$=[Y^{i,j}-h\bar{V}_{i}^{\prime}(Y^{i,j},q_{*}^{-i})]-\int_{0}^{h}[\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})-\bar{V}_{i}^{\prime}(Y_{0}^{i,j},q_{*}^{-i})]\mathrm{d}t+\sqrt{2}B_{h}^{i,j}.$
Then,
$W^{i,j}-Y_{h}^{i,j}$ $=\Big{(}\underbrace{[U^{i,j}-h\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{X}_{n}}^{-i})]}_{Y_{+}^{i,j}}-\underbrace{[Y^{i,j}-h\bar{V}_{i}^{\prime}(Y^{i,j},q_{*}^{-i})]}_{U_{+}^{i,j}}\Big{)}$
$\quad-\underbrace{\int_{0}^{h}[\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})-\bar{V}_{i}^{\prime}(Y_{0}^{i,j},q_{*}^{-i})]\mathrm{d}t}_{A^{i,j}}.$
We have
$\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}=\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]+\mathbb{E}|A^{i,j}|^{2}.$ (5.6)
We make two claims and defer their proofs to Sections 5.4 and 5.5, respectively.
Claim 5.1.
For any $\varepsilon>0$,
$-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]+\mathbb{E}|A^{i,j}|^{2}$
$\leq\varepsilon h\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+(\varepsilon^{-1}+h)h^{3}[L^{2}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+\widetilde{L}^{2}]+L^{2}h^{3}.$
Claim 5.2.
For any $j\in[N]$,
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}\leq(1-\alpha h)^{2}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})$
By (5.6) and Claim 5.1,
$\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}\leq(1+\varepsilon h)\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+(\varepsilon^{-1}+h)h^{3}[L^{2}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+\widetilde{L}^{2}]+L^{2}h^{3}.$
Based on Claim 5.2,
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}$
$\leq(1+\varepsilon h)\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+(\varepsilon^{-1}+h)h^{3}[L^{2}\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}+m\widetilde{L}^{2}]+mL^{2}h^{3}$
$\leq(1+\varepsilon h)(1-\alpha h)^{2}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+h^{3}m[(\varepsilon^{-1}+h)(L^{4}/\alpha+\widetilde{L}^{2})+L^{2}].$
Take $\varepsilon=\alpha$. We have, $\varepsilon^{-1}+h=\alpha^{-1}+2/(\alpha+L)\leq 2\alpha^{-1}$ and thus,
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}\leq(1-\alpha h)\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+mh^{3}[2\alpha^{-1}(L^{4}/\alpha+\widetilde{L}^{2})+L^{2}].$ (5.7)
5.3.5 Step 5: Establishing the recursive inequality
By (5.3), (5.5), (5.7) and $h\leq B\alpha/(4L^{2})$,
$W_{n+1}^{2}\leq\bigg{(}1-\alpha h+\frac{2L^{2}h^{2}}{B}\bigg{)}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+mh^{3}[2\alpha^{-1}(L^{4}/\alpha+\widetilde{L}^{2})+L^{2}]+\frac{2mh^{2}L^{2}/\alpha}{B}$
$\leq\bigg{[}1-\alpha h\bigg{(}1-\frac{2L^{2}h}{B\alpha}\bigg{)}\bigg{]}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+mh^{2}\max\{h,(LB)^{-1}\}\bigg{(}\frac{2(L^{4}/\alpha+\widetilde{L}^{2})}{\alpha}+L^{2}+L^{3}/\alpha\bigg{)}$
$\leq(1-\alpha h/2)\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+4mh^{2}\max\{h,(LB)^{-1}\}(L^{4}/\alpha^{2}+\widetilde{L}^{2}/\alpha).$
By (5.1),
$(1-\alpha h/2)\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})$ $\leq(1-\alpha h/4)^{2}\bigg{(}W_{n}+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}\bigg{)}^{2}$
$\leq\bigg{(}(1-\alpha h/4)W_{n}+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}\bigg{)}^{2}.$
Therefore,
$W_{n+1}^{2}\leq\bigg{(}(1-\alpha h/4)W_{n}+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}\bigg{)}^{2}+4mh^{2}\max\{h,(LB)^{-1}\}(L^{4}/\alpha^{2}+\widetilde{L}^{2}/\alpha).$
The proof is completed by combining the above estimates.
5.4 Proof of Claim 5.1
Part of the proof follows from the idea of Lemma 24 in Durmus and Moulines (2019). Define $V_{i}(\cdot)=\bar{V}_{i}(\cdot,q_{*}^{-i})$. Note that $Y_{t}^{i,j}\sim q_{*}^{i}$ and the density function of $q_{*}^{i}$ is proportional to $e^{-V_{i}(\cdot)}$. Then,
$\mathrm{d}Y_{t}^{i,j}=-V_{i}^{\prime}(Y_{t}^{i,j})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}^{i,j},$
By Ito’s lemma,
$\mathrm{d}V_{i}^{\prime}(Y_{t}^{i,j})$ $=[-V_{i}^{\prime\prime}(Y_{t}^{i,j})V_{i}^{\prime}(Y_{t}^{i,j})+V_{i}^{\prime\prime\prime}(Y_{t}^{i,j})]\mathrm{d}t+\sqrt{2}V_{i}^{\prime\prime}(Y_{t}^{i,j})\mathrm{d}B_{t}^{i,j}.$
As a result,
$\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})-\bar{V}_{i}^{\prime}(Y_{0}^{i,j},q_{*}^{-i})=V_{i}^{\prime}(Y_{t}^{i,j})-V_{i}^{\prime}(Y_{0}^{i,j})$
$=\underbrace{\int_{0}^{t}[-V_{i}^{\prime\prime}(Y_{s}^{i,j})V_{i}^{\prime}(Y_{s}^{i,j})+V_{i}^{\prime\prime\prime}(Y_{s}^{i,j})]\mathrm{d}s}_{\alpha_{t}^{i,j}}+\sqrt{2}\underbrace{\int_{0}^{t}V_{i}^{\prime\prime}(Y_{s}^{i,j})\mathrm{d}B_{s}^{i,j}}_{\beta_{t}^{i,j}}.$
Let $\{\mathcal{F}_{t}\}_{t\geq 0}$ be the filtration generated by the process $\{(Y_{t}^{i,j},B_{t}^{i,j})\}_{t\geq 0}$ and the array $\bm{X}_{n}$. Since $\alpha_{t}^{i,j}$ is $\mathcal{F}_{t}$-adapted and $\beta_{t}^{i,j}$ is a $\mathcal{F}_{t}$-martingale, we have $\mathbb{E}(\beta_{t}^{i,j}|\mathcal{F}_{t})=0$ and $\mathbb{E}(\alpha_{t}^{i,j}\beta_{t}^{i,j})=0$.
Note that $A^{i,j}=\int_{0}^{h}(\alpha_{t}^{i,j}+\sqrt{2}\beta_{t}^{i,j})\mathrm{d}t$ and
$\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]$ $=\mathbb{E}\bigg{(}(U_{+}^{i,j}-Y_{+}^{i,j})\int_{0}^{h}(\alpha_{t}^{i,j}+\sqrt{2}\beta_{t}^{i,j})\mathrm{d}t\bigg{)}$
$=\int_{0}^{h}\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})(\alpha_{t}^{i,j}+\sqrt{2}\beta_{t}^{i,j})]\mathrm{d}t.$
Since $U_{+}^{i,j}-Y_{+}^{i,j}$ is $\mathcal{F}_{0}$-measurable, $\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})\beta_{t}^{i,j}]=0$. We obtain that
$-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]$ $=-2\int_{0}^{h}\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})\alpha_{t}^{i,j}]\mathrm{d}t$
$\leq\int_{0}^{h}\Big{(}\varepsilon\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+\varepsilon^{-1}\mathbb{E}|\alpha_{t}^{i,j}|^{2}\Big{)}\mathrm{d}t$
$=\varepsilon h\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+\varepsilon^{-1}\int_{0}^{h}\mathbb{E}|\alpha_{t}^{i,j}|^{2}\mathrm{d}t,\qquad\forall\varepsilon>0.$
On the other hand, $A^{i,j}=\int_{0}^{h}(\alpha_{t}^{i,j}+\sqrt{2}\beta_{t}^{i,j})\mathrm{d}t$ and implies that
$\mathbb{E}|A^{i,j}|^{2}$ $=h^{2}\mathbb{E}\bigg{(}\frac{1}{h}\int_{0}^{h}(\alpha_{t}^{i,j}+\sqrt{2}\beta_{t}^{i,j})\mathrm{d}t\bigg{)}^{2}\leq h\int_{0}^{h}\mathbb{E}(\alpha_{t}^{i,j}+\sqrt{2}\beta_{t}^{i,j})^{2}\mathrm{d}t$
$=h\int_{0}^{h}(\mathbb{E}|\alpha_{t}^{i,j}|^{2}+2\mathbb{E}|\beta_{t}^{i,j}|^{2})\mathrm{d}t.$
Consequently,
$-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]+\mathbb{E}|A^{i,j}|^{2}$
$\leq\varepsilon h\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+(\varepsilon^{-1}+h)\int_{0}^{h}\mathbb{E}|\alpha_{t}^{i,j}|^{2}\mathrm{d}t+2h\int_{0}^{h}\mathbb{E}|\beta_{t}^{i,j}|^{2}\mathrm{d}t.$ (5.8)
We now work on $\mathbb{E}|\alpha_{t}^{i,j}|^{2}$ and $\mathbb{E}|\beta_{t}^{i,j}|^{2}$. By Ito isometry and Lemma 5.2, we have
$\mathbb{E}|\beta_{t}^{i,j}|^{2}=\mathbb{E}\bigg{|}\int_{0}^{t}V_{i}^{\prime\prime}(Y_{s}^{i,j})\mathrm{d}B_{s}^{i,j}\bigg{|}^{2}=\mathbb{E}\bigg{(}\int_{0}^{t}|V_{i}^{\prime\prime}(Y_{s}^{i,j})|^{2}\mathrm{d}s\bigg{)}\leq L^{2}t$
and $2h\int_{0}^{h}\mathbb{E}|\beta_{t}^{i,j}|^{2}\mathrm{d}t\leq L^{2}h^{3}$. In addition,
$|\alpha_{t}^{i,j}|^{2}$ $\leq\bigg{(}\int_{0}^{t}[L|V_{i}^{\prime}(Y_{s}^{i,j})|+\widetilde{L}]\mathrm{d}s\bigg{)}^{2}=\bigg{(}L\int_{0}^{t}|V_{i}^{\prime}(Y_{s}^{i,j})|\mathrm{d}t+\widetilde{L}t\bigg{)}^{2}$
$\leq 2L^{2}\bigg{(}\int_{0}^{t}|V_{i}^{\prime}(Y_{s}^{i,j})|\mathrm{d}t\bigg{)}^{2}+2\widetilde{L}^{2}t^{2}=2L^{2}t^{2}\bigg{(}\frac{1}{t}\int_{0}^{t}|V_{i}^{\prime}(Y_{s}^{i,j})|\mathrm{d}t\bigg{)}^{2}+2\widetilde{L}^{2}t^{2}$
$\leq 2L^{2}t^{2}\cdot\frac{1}{t}\int_{0}^{t}|V_{i}^{\prime}(Y_{s}^{i,j})|^{2}\mathrm{d}t+2\widetilde{L}^{2}t^{2}.$
Since $Y_{s}^{:,j}\sim q_{*}$, we have
$\mathbb{E}|V_{i}^{\prime}(Y_{s}^{i,j})|^{2}$ $=\mathbb{E}|\bar{V}_{i}^{\prime}(Y_{s}^{i,j},q_{*}^{-i})|^{2}=\mathbb{E}\Big{|}\mathbb{E}_{\bm{y}\sim q_{*}^{-i}}[\partial_{i}V(Y_{s}^{i,j},\bm{y})]\Big{|}^{2}$
$=\mathbb{E}\Big{|}\mathbb{E}[\partial_{i}V(\bm{Y}_{s}^{:,j})|Y_{s}^{i,j}]\Big{|}^{2}\leq\mathbb{E}|\partial_{i}V(\bm{Y}_{s}^{:,j})|^{2}=\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}.$
and
$\mathbb{E}|\alpha_{t}^{i,j}|^{2}\leq 2t^{2}[L^{2}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+\widetilde{L}^{2}].$
Hence,
$\int_{0}^{h}\mathbb{E}|\alpha_{t}^{i,j}|^{2}\mathrm{d}t\leq\frac{2h^{3}}{3}[L^{2}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+\widetilde{L}^{2}].$
Finally, we obtain from (5.8) that
$-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]+\mathbb{E}|A^{i,j}|^{2}$
$\leq\varepsilon h\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+(\varepsilon^{-1}+h)h^{3}[L^{2}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+\widetilde{L}^{2}]+L^{2}h^{3}.$
5.5 Proof of Claim 5.2
Denote by $\mathcal{F}^{i}$ the $\sigma$-field generated by $(\bm{X}_{n},\bm{Y}^{i,:})$. Fix and $(i,j)\in[m]\times[N]$. $\bm{U}^{i,:}$ are $\mathcal{F}^{i}$-measurable. Fact 5.1 implies that conditioned on $\mathcal{F}^{i}$, we have $\bm{Y}^{-i,j}\sim q_{*}^{-i}$ and $\bm{U}^{-i,j}\sim q_{\bm{X}_{n}}^{-i}$. As a result,
$\mathbb{E}[\partial_{i}V(\bm{Y}^{:,j})|\mathcal{F}^{i}]=\mathbb{E}[\partial_{i}V(Y^{i,j},\bm{Y}^{-i,j})|\mathcal{F}^{i}]=\mathbb{E}_{\bm{y}\sim q_{*}^{-i}}\partial_{i}V(Y^{i,j},\bm{y})=\bar{V}_{i}^{\prime}(Y^{i,j},q_{*}^{-i}),$
$Y_{+}^{i,j}=Y^{i,j}-h\bar{V}_{i}^{\prime}(Y^{i,j},q_{*}^{-i})=\mathbb{E}[Y^{i,j}-h\partial_{i}V(\bm{Y}^{:,j})|\mathcal{F}^{i}].$
Similarly, we have $\mathbb{E}[\partial_{i}V(\bm{U}^{:,j})|\mathcal{F}^{i}]=\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{X}_{n}}^{-i})$ and
$U_{+}^{i,j}=U^{i,j}-h\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{X}_{n}}^{-i})=\mathbb{E}[U^{i,j}-h\partial_{i}V(\bm{U}^{:,j})|\mathcal{F}^{i}].$
Consequently,
$\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}$ $=\mathbb{E}\Big{|}\mathbb{E}\Big{(}[U^{i,j}-h\partial_{i}V(\bm{U}^{:,j})]-[Y^{i,j}-h\partial_{i}V(\bm{Y}^{:,j})]\Big{|}\mathcal{F}^{i}\Big{)}\Big{|}^{2}$
$\leq\mathbb{E}\Big{|}[U^{i,j}-h\partial_{i}V(\bm{U}^{:,j})]-[Y^{i,j}-h\partial_{i}V(\bm{Y}^{:,j})]\Big{|}^{2}.$
Define $\bm{\phi}:~{}\mathbb{R}^{mr}\to\mathbb{R}^{mr}$, $\bm{x}\mapsto\bm{x}-h\nabla V(\bm{x})$. Then,
$\sum_{i=1}^{m}\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}\leq\mathbb{E}\|\bm{\phi}(\bm{U}^{:,j})-\bm{\phi}(\bm{Y}^{:,j})\|_{2}^{2}\leq(1-\alpha h)^{2}\mathbb{E}\|\bm{U}^{:,j}-\bm{Y}^{:,j}\|_{2}^{2},$
where we used Lemma 6.2 and the assumption $0<h<2/(\alpha+L)$. Summing over $(i,j)$, we get
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}\leq(1-\alpha h)^{2}\mathbb{E}\|\bm{U}-\bm{Y}\|_{\mathrm{F}}^{2}/N.$
The proof is finished by (5.2).
6 Technical lemmas
Lemma 6.1 (Additivity of $\mathcal{W}_{2}$ for product measures).
Let $m\in\mathbb{Z}_{+}$ and $p_{i},q_{i}\in\mathcal{P}(\mathbb{R})$ for all $i\in[m]$. Then,
$\mathcal{W}_{2}^{2}\bigg{(}p_{1}\otimes\cdots\otimes p_{m},~{}q_{1}\otimes\cdots\otimes q_{m}\bigg{)}=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(p_{i},q_{i}).$
Proof of Lemma 6.1.
For any $i\in[m]$, let $\gamma_{i}$ be a $\mathcal{W}_{2}$-optimal coupling between $p_{i}$ and $q_{i}$. Then, $\gamma_{1}\otimes\cdots\otimes\gamma_{m}$ is a coupling between $p$ and $q$. For $(\bm{X},\bm{Y})\sim\gamma$, we have
$\mathcal{W}_{2}^{2}\bigg{(}p_{1}\otimes\cdots\otimes p_{m},~{}q_{1}\otimes\cdots\otimes q_{m}\bigg{)}\leq\mathbb{E}\|\bm{X}-\bm{Y}\|_{2}^{2}=\sum_{i=1}^{m}\mathbb{E}|X_{i}-Y_{i}|^{2}=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(p_{i},q_{i}).$
To prove the opposite direction of the bound, define $\gamma$ as a $\mathcal{W}_{2}$-optimal coupling between $p$ and $q$. If we draw $(\bm{X},\bm{Y})\sim\gamma$, then $X_{i}\sim p_{i}$, $Y_{i}\sim q_{i}$, and
$\mathcal{W}_{2}^{2}\bigg{(}p_{1}\otimes\cdots\otimes p_{m},~{}q_{1}\otimes\cdots\otimes q_{m}\bigg{)}=\mathbb{E}\|\bm{X}-\bm{Y}\|_{2}^{2}=\sum_{i=1}^{m}\mathbb{E}|X_{i}-Y_{i}|^{2}\geq\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(p_{i},q_{i}).$
Lemma 6.2 (Lemma 1 in Dalalyan (2017a)).
Let Assumption 4.1 hold and $0<h<2/(\alpha+L)$. Define $\bm{\phi}(\bm{x})=\bm{x}-h\nabla V(\bm{x})$, $\forall\bm{x}\in\mathbb{R}^{mr}$. We have
$\|\bm{\phi}(\bm{x})-\bm{\phi}(\bm{y})\|_{2}\leq(1-\alpha h)\cdot\|\bm{x}-\bm{y}\|_{2},\qquad\forall\bm{x},\bm{y}\in\mathbb{R}^{mr}.$
Lemma 6.3.
Under Assumption 4.1, we have $\mathbb{E}_{\bm{x}\sim q_{*}}[\nabla V(\bm{x})]=\bm{0}$, $\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}\leq mL^{2}/\alpha$, and $\operatorname{\rm cov}_{\bm{x}\sim q_{*}}(\bm{x})\preceq\alpha^{-1}\bm{I}_{m}$.
Proof of Lemma 6.3.
By definition,
$\mathbb{E}_{\bm{x}\sim q_{*}}\partial_{1}V(\bm{x})=\mathbb{E}_{x_{1}\sim q_{*}^{1}}[\mathbb{E}_{\bm{x}^{-1}\sim q_{*}^{-1}}\partial_{1}V(x_{1},\bm{x}^{-1})]=\mathbb{E}_{x_{1}\sim q_{*}^{1}}\bar{V}_{1}^{\prime}(x_{1},q_{*}^{-1}).$
Define $V_{1}(\cdot)=\bar{V}_{1}(\cdot,q_{*}^{-1})$. The density function of $q_{*}^{1}$ is proportional to $e^{-V_{1}(\cdot)}$, $V_{1}^{\prime\prime}(\cdot)\geq\alpha$ (Lemma 5.2), and thus $\lim_{x\to\pm\infty}V_{1}(x)=+\infty$. Hence,
$\mathbb{E}_{x_{1}\sim q_{*}^{1}}\bar{V}_{1}^{\prime}(x_{1},q_{*}^{-1})\propto\int_{\mathbb{R}}V_{1}^{\prime}(x)e^{-V_{1}(x)}\mathrm{d}x=-\int_{\mathbb{R}}\mathrm{d}e^{-V_{1}(x)}=0.$
Based on the above, $\mathbb{E}_{\bm{x}\sim q_{*}}\nabla V(\bm{x})=\bm{0}$.
Consequently, $\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{1}V(\bm{x})|^{2}=\operatorname{{\rm var}}_{\bm{x}\sim q_{*}}[\partial_{1}V(\bm{x})]$. Note that $q_{*}$ is a product distribution with strongly log-concave marginals (Lemma 5.2). According to Bakry and Émery (2006), $q_{*}$ satisfies the Poincaré inequality
$\operatorname{{\rm var}}_{\bm{x}\sim q_{*}}f(\bm{x})\leq\alpha^{-1}\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla f(\bm{x})\|_{2}^{2},\qquad\forall f\in C^{1}(\mathbb{R}^{m}).$ (6.1)
Hence,
$\operatorname{{\rm var}}_{\bm{x}\sim q_{*}}[\partial_{1}V(\bm{x})]\leq\alpha^{-1}\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla[\partial_{1}V(\bm{x})]\|_{2}^{2}=\alpha^{-1}\sum_{j=1}^{m}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial^{2}_{j1}V(\bm{x})|^{2}.$
By Assumption 4.1,
$\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}$ $=\alpha^{-1}\sum_{i,j\in[m]}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial^{2}_{ij}V(\bm{x})|^{2}=\alpha^{-1}\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla^{2}V(\bm{x})\|_{\mathrm{F}}^{2}$
$\leq\alpha^{-1}\mathbb{E}_{\bm{x}\sim q_{*}}[m\|\nabla^{2}V(\bm{x})\|_{2}^{2}]\leq mL^{2}/\alpha.$
Finally, choose any unit vector $\bm{u}\in\mathbb{R}^{m}$. By (6.1), we have
$\operatorname{{\rm var}}_{\bm{x}\sim q_{*}}(\langle\bm{u},\bm{x}\rangle)\leq\alpha^{-1}\mathbb{E}_{\bm{x}\sim q_{*}}\|\bm{u}\|_{2}^{2}=\alpha^{-1}.$ (6.2)
Therefore, $\operatorname{\rm cov}_{x\sim q_{*}^{i}}(\bm{x})\leq\alpha^{-1}\bm{I}_{m}$. ∎
Lemma 6.4 (Corollary 6.12 in Bobkov and Ledoux (2019)).
Let $\mu\in\mathcal{P}(\mathbb{R})$ be log-concave with variance $\sigma^{2}$, and $\{X_{j}\}_{j=1}^{N}$ be i.i.d. samples from $\mu$. There exists a universal constant $c$ such that
$\mathbb{E}\mathcal{W}_{2}^{2}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\delta_{X_{i}},\mu\bigg{)}\leq\frac{c\sigma^{2}\log N}{N},\qquad\forall N\geq 2.$
Corollary 6.1.
Let Assumption 4.1 hold and $N\geq 2$. Consider the notations in Definition 3.2, and define $\bm{Y}\in\mathbb{R}^{m\times N}$ with $\{\bm{Y}^{:,j}\}_{j=1}^{N}$ being i.i.d. samples from $q_{*}$. There exists a universal constant $c$ such that
$\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})\leq\frac{cm\log N}{\alpha N}.$
Proof of Corollary 6.1.
By Lemma 6.1,
$\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(q_{\bm{Y}}^{i},q_{*}^{i})=\mathcal{W}_{2}^{2}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\delta_{Y^{i,j}},q_{*}^{i}\bigg{)}.$
By Lemmas 6.3 and 6.4, there exists a universal constant $c$ such that
$\mathbb{E}\mathcal{W}_{2}^{2}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\delta_{Y^{i,j}},q_{*}^{i}\bigg{)}\leq\frac{c\log N}{\alpha N},\qquad\forall i\in[m],$
and thus,
$\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})\leq\frac{cm\log N}{\alpha N}.$
7 Discussions
In this paper, we introduce PArticle VI (PAVI), a particle algorithm for mean-field variational inference (MFVI). The algorithm is based on formulating the relevant optimization problem in terms of Fokker-Planck equations and corresponding diffusion processes. We also provide a theoretical analysis of the particle algorithm, which provides a non-asymptotic upper bound on the $2$-Wasserstein distance between the empirical measure of the particles at the $n$th iteration and the solution of the MFVI optimization problem. Several future directions are worth pursuing. One may consider relaxing the regularity assumptions in Theorem 4.1 and adapting PAVI to the setting of parametric MFVI. Another interesting direction is to develop accelerated versions of PAVI using the recipe in the recent works Wang and Li (2022); Chen et al. (2023).
Acknowledgement
The authors thank Daniel Lacker for helpful discussions. Du’s research is supported in part by National Science Foundation grants DMS-2309245 and DMS-1937254, and the Department of Energy grant DE-SC0025347. Wang’s research is supported by the National Science Foundation grant DMS-2210907.